OpenStudy (anonymous):

Solve the equation (10-6x)^1/2=(3x+4)^1/2

OpenStudy (anonymous):

square both sides 10-6x = 3x+4 Isolate x 9x = 6 divide by 9 x = 1/3

OpenStudy (anonymous):

That is not one of the choices

OpenStudy (anonymous):

that's because I can't divide :). It should be 2/3

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem